OpenStudy (ny,ny):

The line that passes through the points (-5,2) and (1,y) has a slope of 5/6. Find y.

OpenStudy (lgbasallote):

the slope of the two coordinates must be equal to 5/6 therefore \[\frac{y_2 -y_1}{x_2 -x_1} = \frac{5}{6}\] \[\frac{y-2}{1-(-5)} = \frac{5}{6}\] \[\frac{y - 2}{6} = \frac{5}{6}\] 6 can be cancelled... \[y - 2 = 5\] gpt it?

OpenStudy (ny,ny):

yeahhh. so y=7.

OpenStudy (lgbasallote):

yup!

OpenStudy (ny,ny):

cool. thank yu =)

OpenStudy (ny,ny):

sorry for taking so long.

OpenStudy (lgbasallote):

nahh no biggie ^_^